### v2

# CMQCIC-Bench: A Chinese Benchmark for Evaluating Large Language Models in Medical Quality Control Indicator Calculation

Abstract
Medical quality control indicators are essential to assess the qualifications of healthcare institutions for medical services. With the impressive performance of large language models (LLMs) like GPT-4 in the medical field, leveraging these technologies for the Medical Quality Control Indicator Calculation (MQCIC) presents a promising approach. In this work, (1) we introduce a real-world task MQCIC and propose an open-source Chinese electronic medical records (EMRs)-based dataset (CMQCIC-Bench) comprising 785 instances and 76 indicators. (2) We propose a semi-automatic method to enhance the rule representation. Then we propose the Clinical Facts-based Inferential Rule (CF-IR) method that disentangles the clinical fact verification and inferential rule reasoning actions. (3) We conduct comprehensive experiments on 20 representative LLMs, covering general and medical models. Our findings reveal that CF-IR outperforms Chain-of-Thought methods in MQCIC tasks. (4) We conduct an error analysis and investigate the capabilities of clinical fact verification and inferential rule reasoning, providing insights to improve performance in the MQCIC further. The dataset and code is available in this repository 111https://github.com/YuY-2001/C-MQCIC.
[…]
CMQCIC-Bench: A Chinese Benchmark for Evaluating Large Language Models in Medical Quality Control Indicator Calculation
Guangya Yu, Yanhao Li, Zongying Jiang, Yuxiong Jin, Li Dai, Yupian Lin, Ruihui Hou, Weiyan Zhang, Yongqi Fan, Qi Ye††thanks: Corresponding authors., Jingping Liu, Tong Ruan11footnotemark: 1 School of Information Science and Engineering, East China University of Science and Technology, Shanghai, China guangyayu@mail.ecust.edu.cn, {yeh_qi1125,ruantong}@ecust.edu.cn
1 Introduction
Medical quality control indicators play an essential role in assessing the performance of healthcare institutions Øvretveit (2001); Wang et al. (2018); Anderson et al. (2017). Recently, Large Language Models (LLMs) like GPT-4 Achiam et al. (2023) have shown promising capabilities in the medical domain. These include applications such as diagnostic reasoning Dou et al. (2024), clinical note generation Yang et al. (2023a), and automated clinical assessment GU et al. (2024). Such capabilities also prove effective in Medical Quality Control Indicator Calculation (MQCIC) Ye et al. (2025).
Traditionally, calculating quality indicators relied on manually constructed rules (regular expressions) Tamang et al. (2015); Hsu et al. (2016), which is time-consuming Ross et al. (2015). As illustrated in Figure 1, the indicator "ICH Improvement/Stabilization at Discharge Rate" contains <Definition, Calculation Formula>, which requires (i) medical knowledge regarding the Glasgow Coma Score(GCS) and NIH Stroke Score(NIHSS), reflecting different patient conditions; (ii) multiple sources of information, including both admission and discharge records; and (iii) numerical computation or logical reasoning. With such fine-grained rules, experts develop scripts to identify the relevant data from the unstructed text. However, these quality control indicators related to various diseases are continually refined and expanded over time. Relying solely on fixed scripts or NLP extraction methods is inadequate and lacks generalizability Lee et al. (2019); Raju et al. (2015).
Fortunately, LLMs have demonstrated excellent performance in the transformation as well as decomposition of rules Wang et al. (2023a); Wu et al. (2024); Xu et al. (2024). However, several obstacles remain in developing LLM-based clinical applications Huang et al. (2024), especially MQCIC: (i) LLMs struggle to provide accurate, reliable answers for complex clinical reasoning tasks, especially when using Chain-of-Thought (CoT) reasoning Wei et al. (2022). (ii) Concerns over LLMs’ reliance on opaque, "black-box" methods for clinical decisions, which may erode user trust. Increasing focus is being placed on improving LLM reasoning to follow explicit logical rules Servantez et al. (2024a); Yang et al. (2024c); Sun et al. (2024), shifting from context-dependent to transparent, rule-based prompting.
Therefore, this work aimed to explore leveraging explicit rules to achieve automated indicator calculation using LLMs based on electronic medical records (EMRs). Firstly, we introduce a real-world task MQCIC, and propose an open-source dataset, CMQCIC-Bench, derived from Chinese EMRs on an online Chinese website. The dataset comprises 785 instances spanning 76 indicators. Each instance consists of a Patient Note, a Question, and an Answer. We also provide detailed annotations of clinical facts and explanations. Due to the ambiguity of existing rules that impairs the effectiveness of LLMs, we propose a semi-automatic method to enhance the rule representation. With these refined rules, we introduce the Clinical Fact-based Inferential Rule reasoning (CF-IR) method that disentangles the two abilities during the inference stage. We conducted extensive experiments on 20 representative LLMs across general and medical domain. The evaluation results demonstrate that CF-IR outperforms the CoT method. Furthermore, we investigated the capabilities of clinical fact verification and inferential rule reasoning.
In summary, the major contributions are as follows:
We introduce a clinical scenario task Medical Quality Control Indicator Calculation and propose CMQCIC-Bench, a Chinese open-source dataset with 785 instances, covering 76 different medical quality control indicators.
We propose a semi-automatic method to enhance the rule representation. Then we propose the Clinical Fact-based Inferential Rule reasoning (CF-IR) method that disentangles the clinical fact verification and inferential rule reasoning actions.
We conducted comprehensive experiments on 20 representative LLMs, where CF-IR improved performance by 0.43% in the zero-shot setting and 1.45% in the one-shot setting.
We analyze errors and explore clinical fact verification and rule reasoning, offering insights to improve MQCIC performance.
2 The Medical Quality Control Indicator Calculation Task
Typically, MQCIC involves three steps: (1) Recall relevant EMRs from all cases based on the denominator rules of the indicator. (2) Identify the EMRs that meet the numerator rules from these relevant EMRs. (3) Finally, compute the proportion to determine the indicator’s value. The first step can be addressed by matching the ICD-10 codes with diagnostic results. However, the second step is the most challenging, which is the focus of this work. Considering the type of answer is not unique, we define the task as a binary classification problem rather than a cloze task. Thus, the problem is defined as follows: given a Patient Note and a Question related to the indicator’s rule, the task of MQCIC is to generate the answer .
3 Dataset Construction
In this section, we construct a dataset, CMQCIC-Bench, for the MQCIC task. The main content includes the data collection of indicators and patient notes, data annotation, and data characteristics.
3.1 Data Collection
We collected indicators and patient notes from two sources. Indicators Sources. We manually curated 76 challenging indicators from authoritative documents222http://www.ncis.cn/home, all developed by experts. For each indicator, a rule-related question was constructed for inclusion in the CMQCIC-Bench. Patient Notes Sources. We gathered raw data from a Chinese open-source medical website333https://www.iiyi.com/. Patient notes meeting the denominator rules were filtered based on ICD-10 codes and diagnostic findings. Finally, we manually removed patient names, hospital information, and other sensitive data to ensure no privacy risks.
3.2 Data Annotation
Specifically, the annotation process uses the following three-step pipeline. (1) Clinical fact extraction. We leverage GPT-4o to extract the original information from EMRs without any modification, then reason based on the context to verify the clinical fact. The clinical facts contain GCS scores, lab exams, medications, etc. (2) Answer and explanation generation. Given the obtained facts, for each instance, we leverage GPT-4o to generate the step-by-step explanation through logical reasoning and a final answer . (3) Data quality control. Finally, with the guidance of medical experts, annotators are required to check the answer in three facets: fact extraction, logical reasoning, and consistency. Fact extraction and logical reasoning verify the accuracy of the first two stages, while consistency ensures alignment within the patient notes to exclude low-quality cases. In the end, we curated 785 instances for CMQCIC-Bench, as shown in Figure 2, each instance consists of a Patient Note , a Question , a step-by-step , and the final answer . With the same process, we constructed a CMQCIC-Private dataset derived from patient notes of top-tier tertiary hospitals in China. Ethics committees and experts have rigorously de-identified these data to ensure no privacy leakage risk.
3.3 Data Characteristics
As shown in Table 1, we use Tiktoken444https://github.com/openai/tiktoken to measure sample lengths, yielding average lengths of 380.41 and 520.71, respectively. The shorter average length of each in CMQCIC-Bench compared to the private dataset stems from the summarized nature of the source data (e.g., lab exams include only key findings). Despite this, the number of facts ranges from 1 to 13, with averages of 3.59 and 4.02, underscoring the task’s demand for multi-step reasoning, consistent with real-world scenarios. Additionally, Figure 3 illustrates the indicator distribution, which spans 30 diseases.
| CMQCIC-Bench | CMQCIC-Private | |
| Indicators | 76 | 42 |
| Instance | 785 | 314 |
| Avg. L of Note | 380.41 | 520.71 |
| Avg. L of Q. | 99.72 | 113.04 |
| Min Facts | 1 | 1 |
| Max Facts | 13 | 13 |
| Avg. Facts | 3.59 | 4.02 |
| Open-source | Yes | No |
4 Method
As shown in Figure 1, the indicators constructed by experts are quite vague, and the underlying medical knowledge can directly affect the implementation of the rules. Therefore, we propose a semi-automatic method that decomposes them into transparent, templated clinical facts and logical rules. We then introduce the Clinical Fact-based Inferential Rule (CF-IR) reasoning pattern for inference. As illustrated in Figure 4, the method comprises two key components: Rule Representation Enhancement and CF-IR. Additionally, we introduce Automatic CF-IR (ACF-IR) to explore the automatic performance of rule representation enhancement.
4.1 Rule Representation Enhancement
Step 1: Knowledge Enhancement. We leverage GPT-4o to recall relevant information instead of collecting additional medical books and guidelines. It aims to resolve rule ambiguities arising from the lack of knowledge. For example, the high and low scores of GCS and NIHSS have different meanings, and splitting the rules based solely on semantic information could result in incorrect logical rules.
Step 2: Rule Decomposition. We use GPT-4o to break down complex rules into simpler logical rules in both natural and symbolic language. Symbolic language streamlines natural language by converting it into variables, combined with mathematical and logical symbols to build logical expressions.
Step 3: Clinical Fact Templatization. GPT-4o further extracts the clinical facts involved in the logical rules, which are independent. Each clinical fact with a supposed answer set, such as True/False, numerical value unit, etc.
In the end, we enlisted human experts to review the enhanced rules for the 76 indicators.
4.2 Clinical Fact-based Inferential Rule Reasoning
Motivated by Jin et al. (2024), the model performs two steps during inference: Clinical Fact Verification and Inferential Rule Reasoning. LLM first extracts and verifies the related information using the templated clinical fact. Then, LLM reasoning on the verified clinical facts with explicit logical rules. We believe that disentangling these two distinct abilities improves performance during the inference process and enhances interpretability. Here are some discussions about the two abilities, and we explore deeply in Section 6.3.
(1) Clinical Fact Verification. Before engaging in clinical reasoning, it is crucial to obtain accurate clinical informationWang et al. (2024a). However, extracting clinical facts and verifying them through reasoning in a long context with noises of over 380.41 tokens is quite challenging. This includes identifying synonyms, linking symptoms to facts (e.g., high GCS indicates better consciousness), understanding medications (e.g., dual antiplatelet therapy), and analyzing surgical indications, all requiring medical knowledge and clinical reasoning.
| Zero-Shot | One-Shot | ||||||
| Standard | CoT | CF-IR | ACF-IR | CoT | CF-IR | ||
| General | MiniCPM3-4B Hu et al. (2024) | 63.31 | 72.10 | 68.91 | 78.98 | 83.56 | 82.67 |
| Internlm2.5-1.8B Cai et al. (2024) | 56.17 | 56.18 | 54.14 | 65.85 | 68.91 | 64.07 | |
| Internlm2.5-7B Cai et al. (2024) | 63.31 | 73.12 | 79.49 | 77.07 | 84.07 | 84.45 | |
| Internlm2.5-20B Cai et al. (2024) | 69.04 | 77.57 | 80.63 | 83.06 | 86.36 | 88.78 | |
| Qwen2.5-0.5B Yang et al. (2024a) | 54.26 | 56.05 | 53.88 | 52.86 | 61.01 | 53.63 | |
| Qwen2.5-1.5B Yang et al. (2024a) | 66.11 | 63.43 | 62.42 | 60.50 | 71.21 | 73.24 | |
| Qwen2.5-3B Yang et al. (2024a) | 60.38 | 73.37 | 67.64 | 79.36 | 77.83 | 82.03 | |
| Qwen2.5-7B Yang et al. (2024a) | 66.49 | 82.80 | 82.92 | 85.73 | 85.22 | 89.93 | |
| Qwen2.5-14B Yang et al. (2024a) | 78.98 | 82.03 | 86.11 | 84.96 | 87.89 | 91.59 | |
| Qwen2.5-32B Yang et al. (2024a) | 75.54 | 86.49 | 87.21 | 92.35 | 89.80 | 94.77 | |
| Qwen2.5-72B Yang et al. (2024a) | 87.77 | 87.51 | 92.73 | 91.33 | 90.95 | 95.54 | |
| llama3.1-8B Dubey et al. (2024) | 48.53 | 63.05 | 78.34 | 73.88 | 81.52 | 85.85 | |
| llama3.1-70B Dubey et al. (2024) | 82.54 | 85.85 | 85.47 | 84.45 | 88.53 | 91.84 | |
| GPT-4o Achiam et al. (2023) | 77.45 | 88.91 | 91.84 | 90.57 | 91.59 | 93.88 | |
| Medical | HuatuoGPT2-7B Chen et al. (2024) | 54.01 | 54.26 | 49.55 | 48.66 | 53.50 | 56.81 |
| HuatuoGPT2-14B Chen et al. (2024) | 53.63 | 55.28 | 46.36 | 37.19 | 52.10 | 43.31 | |
| Apollo2-0.5B Zheng et al. (2025) | 39.55 | 41.96 | 41.14 | 57.19 | 54.39 | 65.47 | |
| Apollo2-1.5B Zheng et al. (2025) | 53.31 | 52.03 | 50.82 | 52.61 | 66.11 | 65.22 | |
| Apollo2-7B Zheng et al. (2025) | 57.57 | 60.00 | 61.91 | 48.91 | 71.71 | 65.35 | |
| Apollo-72B Wang et al. (2024c) | 68.91 | 76.24 | 72.61 | 80.63 | 86.11 | 86.36 | |
| Average | 63.84 | 69.41 | 69.71 | 71.31 | 76.62 | 77.73 | |
| Human | 95.00 | ||||||
(2) Inferential Rule Reasoning. Reasoning can be categorized in multiple ways Yu et al. (2024). In this study, we specifically define ’Inferential Rule Reasoning’ as the capability to deduce a final conclusion by applying logical rules to multiple clinical facts. Specifically, for each templated clinical fact, the LLM first performs targeted information extraction and verification. Subsequently, it applies explicit logical rules to conduct structured reasoning based on the verified clinical facts.
5 Experiments
We now describe the experimental setup, empirically evaluate CF-IR, and compare it with existing methods using the CMQCIC-Bench dataset.
5.1 Settings
Here are the details of the models, methods, and evaluation metrics.
Models. We select a representative set of 14 general large language models including GPT-4555gpt-4o-2024-08-06 Achiam et al. (2023), Qwen2.5 series Yang et al. (2024a), Internlm2.5 series Cai et al. (2024), Llama3.1 series Dubey et al. (2024), MiniCPM3-4B Hu et al. (2024), as well as 6 medical large language models, comprising the Apollo series Zheng et al. (2025); Wang et al. (2024c), and HuatuoGPT2 series Chen et al. (2024). We omit the "chat" and "instruct" in Table 2.
| Type | Question | Patient Note | Error of LLMs |
| A | Can the screening for anxiety and depression symptoms be considered complete based on the medical records of hospitalized Parkinson’s disease patients? | On the initial visit and the 34th day after treatment, the patient’s H-Y stage decreased from stage 3 to 2.5, the HAMD score dropped from 23 to 2, the Beck Depression Inventory (BDI) score decreased from 14 to 4, the PDSS-2 score dropped from 10 to 2,[…]. | Based on the provided electronic medical record information, there are no clear mentions of specific results for anxiety and depression symptom screenings. Therefore, we are unable to determine the values for the following facts: Anxiety symptom screening = Not sure Depression symptom screening = Not sure. ##HAMD and BDI can evaluate the depression. |
| B | Please assess if the patient’s coronary intervention meets post-operative success criteria: residual stenosis <20% after stenting or <50% after PTCA, with a TIMI grade of 3. | Treatment Process: […] Post-stent implantation, there was no significant residual stenosis at the LAD lesion and TIMI grade II. | Based on the provided instructions and medical records, […] After stent implantation, there was no significant residual stenosis in the patient’s left coronary artery LAD. Therefore, the answer is "Yes" ##comment: TIMI grade must be III. |
Methods. We assess six prompt-based methodologies: (1) Standard Prompt, utilizing solely the original rules and patient notes; (2) Zero-Shot CoT Kojima et al. (2022), enhanced with the directive "Let’s think step by step"; (3) For CF-IR prompt, we leverage human-reviewd fact templates and logical rules to derive the answer; (4) Specifically, we examine the One-shot CoT Wei et al. (2022); (5) To explore the ability of LLM for rule representation enhancement, we introduce ACF-IR, an automated framework that enables LLMs to decompose rules and then conduct CF-IR; (6) Additonally, we set One-Shot CF-IR.
For each indicator, we selected an example outside of CMQCIC-Bench. The outputs for these examples were generated by GPT-4o in a zero-shot setting and were carefully reviewed and annotated by human evaluators.
Evaluation. Following prior works Servantez et al. (2024b); Yang et al. (2024c), the final answer for each instance was labeled as ’True/Yes’ or ’False/No’, enabling us to use accuracy as the outcome evaluation metric. To evaluate step-wise reasoning, we compared responses against ground truth using DeepSeek Liu et al. (2023), assessing two dimensions: Fact Faithfulness (relevance to the original text) and Fact Correctness (accuracy of the fact result). Scores range from 0 (irrelevant/incorrect) to 1 (fully relevant/correct). We utilized DeepSeek 666https://www.deepseek.com/ to extract facts from gold explanations and score model responses accordingly. Scores were averaged per instance to accommodate varying fact counts, resulting in an overall average score. Below is a formal definition of the metrics:
where Judge() represents the LLMs, outputting 0 or 1. The m denotes the number of facts in the i-th instance. The denotes the j-th fact of the i-th instance. Human evaluation, we designed regex to extract key information, subsequently assessed by experts.
Implement Details. We conduct all experiments on H800 and use VLLM 777https://github.com/vllm-project/vllm to accelerate for general LLMs. Specifically, we load the medical LLMs directly. Additionally, we set the max_new_tokens = 1024; repetition_penalty = 1.2; temperature = 0.001. The experiments were run three times with random seeds, and the scores were averaged.
5.2 Main Results
Table 2 presents our evaluation results of various LLMs on the CMQCIC-Bench dataset.
(1) Current leading general LLMs perform better than medical LLMs. Qwen2.5-32B/72B-Instruct, and GPT-4o score similarly at 94.77, 95.54, and 93.88, respectively, while medical LLMs lag, with Apollo-72b scoring only 86.36. Only Qwen2.5-72B-Instruct nears human performance, highlighting the ongoing challenge of the MQCIC task for current methods and LLMs.
(2) CF-IR methods perform better than CoT across different parameters and models. In zero-shot and one-shot settings, the average score of CF-IR improves by 0.43% and 1.45%, respectively, compared to CoT. Unlike the CoT method, which performs reasoning along random paths, our approach integrates explicit logical rules with verifiable facts, enhancing the stability and interpretability of LLMs. While CF-IR demonstrated strong performance across various parameters in the one-shot setting, we observed that in the zero-shot scenario, CF-IR outperformed CoT only on general models with parameters 7B. We will analyze our improvement in Section 6.
(3) One-Shot setting can bring significant improvements. In general, after providing the examples, CoT and CF-IR achieved improvements of 10.38% and 11.50%, respectively, the performance of all models showed significant improvements in the one-shot setting for both the CoT and CF-IR methods except HuatuoGPT2. This may stem from HuatuoGPT’s fine-tuning data being predominantly centered around QA tasks Chen et al. (2024), without incorporating clinical scenarios, and weak in instruction-following.
(4) Automated rule representation enhancement remains challenging. While CF-IR achieves strong performance (77.73) with enhanced rule representation, ACF-IR’s automated approach scores lower (71.31), underperforming CoT. Notably, only Apollo2-0.5B and specific Qwen2.5 variants (3B, 7B, 32B, 72B) surpass CoT in one-shot settings, revealing the limitations of intrinsic model planning capabilities Servantez et al. (2024b); Yang et al. (2024b). A promising direction involves leveraging advanced open-source models (e.g., GPT-4o) or specialized plan training Wu et al. (2024) for rule decomposition, complemented by medical models for inference.
6 Empirical Analysis and Discussion
In this section, we analyze errors and evaluate step-wise reasoning, further exploring clinical fact verification and inferential rule reasoning capabilities.
6.1 Error Analysis
Firstly, we categorize errors into three types: Type A, B and C, representing errors in clinical fact verification, inferential rule reasoning, and other types, respectively. We display the example of two main error types in CMQCIC-Bench in Table 3.
Building on prior work Khandekar et al. (2025), we employ DeepSeek to classify error types by comparing LLM outputs with ground truth in CMQCIC-Bench, facilitating a granular error analysis across LLMs. Since incorrect clinical facts can propagate and affect inferential rule reasoning, we focus on identifying the earliest error type. A manual review of 200 randomly sampled DeepSeek-annotated errors confirmed an 87% accuracy, validating our approach for analyzing error types in all CF-IR responses. As shown in Table 4, providing demonstrations reduces Type A and B errors, highlighting the value of exemplars. While CF-IR does not mitigate clinical fact verification errors, it significantly improves reasoning accuracy due to its structured logical framework. Further details are available in Appendix E.1 (Tables 11 and 12).
| Zero-Shot | One-Shot | |||
| Error Type | CoT | CF-IR | CoT | CF-IR |
| clinical fact | 0.19 | 0.23 | 0.17 | 0.17 |
| reasoning | 0.11 | 0.07 | 0.07 | 0.05 |
| other | 0.00 | 0.01 | 0.00- | 0.00 |
| Total | 0.31 | 0.30 | 0.23 | 0.22 |
6.2 Evaluation on Step-Wise Reasoning
As shown in Table 5, the step-aware evaluation metrics decreased by 14.38 and 17.19 points, respectively, compared to the outcome evaluation results. This suggests that the model often makes clinical fact verification errors during the reasoning process, even when the final result is correct.
| Models | Methods | FC | FF | ACC |
| Qwen2.5-72b | zero-shot CoT | 68.09 | 68.07 | 87.51 |
| zero-shot CF-IR | 76.34 | 76.83 | 92.73 | |
| one-shot CoT | 69.26 | 69.92 | 90.95 | |
| one-shot CF-IR | 90.45 | 86.20 | 95.54 | |
| Qwen2.5-32b | zero-shot CoT | 66.63 | 66.42 | 86.49 |
| zero-shot CF-IR | 72.86 | 71.53 | 87.21 | |
| one-shot CoT | 69.73 | 70.44 | 89.80 | |
| one-shot CF-IR | 84.61 | 76.78 | 94.77 | |
| Qwen2.5-14b | zero-shot CoT | 68.25 | 66.44 | 82.03 |
| zero-shot CF-IR | 70.87 | 65.14 | 86.11 | |
| one-shot CoT | 71.32 | 69.41 | 87.89 | |
| one-shot CF-IR | 83.48 | 78.22 | 91.59 | |
| Qwen2.5-7b | zero-shot CoT | 67.31 | 63.36 | 82.80 |
| zero-shot CF-IR | 67.03 | 65.76 | 82.92 | |
| one-shot CoT | 66.60 | 66.45 | 85.22 | |
| one-shot CF-IR | 77.97 | 71.88 | 89.93 | |
| llama3.1-70b | zero-shot CoT | 65.24 | 62.49 | 85.85 |
| zero-shot CF-IR | 71.70 | 65.23 | 85.47 | |
| one-shot CoT | 69.41 | 68.59 | 88.53 | |
| one-shot CF-IR | 83.15 | 77.50 | 91.84 | |
| llama3.1-8b | zero-shot CoT | 57.02 | 57.52 | 63.05 |
| zero-shot CF-IR | 64.89 | 61.25 | 78.34 | |
| one-shot CoT | 68.13 | 65.61 | 81.52 | |
| one-shot CF-IR | 78.40 | 70.38 | 85.85 | |
| Average | 72.03 | 69.22 | 86.41 | |
6.3 Analysis on Clinical Fact Verification and Inferential Rule Abilities
While the CF-IR method enhances inference performance, we further investigate its two core capabilities: Clinical Fact Verification and Inferential Rule Reasoning. For Clinical Fact Verification, we define the input as <Patient Note, Templated Clinical Fact, Question> and the output as <Reasoning, Final Answer>, evaluated using Fact Faithfulness and Fact Correctness. Unlike step-wise reasoning, we test each fact independently to avoid contextual interference. For Inferential Rule Reasoning, providing verified facts as input to minimize errors, the input is <Verified Clinical Facts, Logical Rules>, and the output is <Explanation, Final Answer>, evaluated using labels like ’True/Yes’ or ’False/No’ for both natural (NL) and symbolic (SY) languages. All experiments are conducted in a zero-shot setting. Additional results are available in Table 13 in Appendix E.2.
As shown in Figures 5, we find that: (1) Both Qwen and Apollo exhibit performance-scale correlations across capabilities. (2) Fact verification performance significantly declines, consistent with Table 5. (3) For inferential reasoning, Qwen performs comparably in natural and symbolic settings, while Apollo shows stronger natural language robustness. (4) With correct facts, Qwen surpasses previous best results (standard, CoT, CF-IR) in zero-shot settings, whereas Apollo underperforms, likely due to Qwen’s extensive logical reasoning training. See Appendix E.2 for additional results.
6.4 Benefit of fine-tuning
We fine-tuned the Qwen2.5-3B-Instruct model using LoRA Hu et al. (2021) for 3 epochs on the CMQCIC-Bench, with evaluation on CMQCIC-Private. The data format follows ACF-IR: Input: <Instruction, Patient Note, Rule>; Output: <Knowledge, Templated Clinical Facts, Logical Rules, Clinical Fact Verification, Inferential Rule Reasoning, Final Answer>. As depicted in Figure 6, the fine-tuned 3B model achieves comparable or superior performance to the 7B model in real-world scenarios. In zero-shot settings, it demonstrates significant improvements across all methods, with gains of 6.7, 8.3, 14.3, and 13.7, confirming the feasibility of distillation rule enhancement for smaller models. More details are provided in Appendix B.
7 Related Work
7.1 Rule-based LLM reasoning
While reasoning demonstrated a fundamental capability of LLM on applications Li et al. (2024), there are many research such as CoT Wei et al. (2022), CoT-sc Wang et al. (2023b), ToT Yao et al. (2024), etc. However, there are more attention on rule-enhanced methods Yang et al. (2023b); Sun et al. (2023); Mu et al. (2023). Reasoning based on facts and deriving answers from logical rules is referred to as inferential rule following ability Sun et al. (2024). Leveraging such ability that integrating explicit rules with LLMs has gained significant attention. For instance, Servantez et al. (2024b) utilized the IRAC framework to tackle legal tasks with LLMs, emphasizing the application of legal rules. Additionally, Wang et al. (2024b) proposed a neurosymbolic framework for multi-step rule application. Despite the current limitations of LLMs in rule-based reasoning Yang et al. (2024c), our work demonstrates that such rule-based reasoning outperforms CoT reasoning in the MQCIC task.
7.2 LLM Evaluations in Clinical Scenarios
While LLMs have shown impressive capabilities in medical knowledge recall and reading comprehension on medical exams Nori et al. (2023); Subramanian et al. (2024), their effectiveness in real-world clinical applications remains a critical area of evaluation. For example, Ouyang et al. (2024) assesses LLMs across 14 expert-curated clinical scenarios, including diagnosis, discharge summaries, and medical consultations. Similarly, Khandekar et al. (2025) introduces MedCal-Bench, a benchmark designed to evaluate inferential rule reasoning in medical contexts, while Hou et al. (2024) simulates a multi-step diagnostic process to test clinical reasoning capabilities. Furthermore, Munnangi et al. (2024); Chung et al. (2025) explore LLMs’ abilities in clinical fact decomposition and verification. In this work, we focus on evaluating LLMs in the MQCIC task, with a specific emphasis on their performance in clinical fact verification and inferential rule reasoning, providing a detailed analysis of these two critical abilities.
8 Conclusion
In this work, we present MQCIC, a novel task, and CMQCIC-Bench, an open-source dataset derived from Chinese EMRs. We propose a semi-automatic approach to refine rule representation and introduce CF-IR, a disentangled inference method. Experimental results show that CF-IR surpasses CoT in performance. Error analysis reveals enhanced capabilities in clinical fact verification and inferential rule reasoning. Additionally, we evaluate step-wise reasoning and conduct a detailed investigation of the two abilities. Our work aims to advance the application of LLMs in MQCIC tasks and offers deeper insights into these essential capabilities.
Limitations and Future Work
While we construct a CMQCIC-Bench dataset and evaluate LLMs’ clinical fact verification and inferential rule reasoning abilities, several limitations can be improved. (1) Due to the difficulty of manually verifying each sample, our dataset only contains 785 instances. (2) We have only located a comprehensive Chinese document on medical quality control indicators. As a result, our dataset consists solely of Chinese EMRs, and we are also leaning toward selecting Chinese LLMs for our analysis. (3) While we observed a significant improvement in model performance with the one-shot demonstration, benchmarking the model with few-shot instances could have further enhanced accuracy, a scenario we did not test. (4) Although we propose the CF-IR method, which performs well across various LLMs with an enhanced rule representation reviewed by humans, decomposing the rules with a smaller LLM that lacks strong planning capabilities remains a challenge.
Ethical Consideration
The medical cases are sourced from the iiyi website, where doctors voluntarily contribute and share their information. The data is explicitly authorized for use in research and educational activities. To safeguard patient privacy, our dataset excludes any personally identifiable details, such as patient names, hospital information, or other sensitive data. As a result, there is no risk of privacy violations related to our dataset. Furthermore, all data usage adheres to ethical guidelines and regulations governing medical information and research.
Acknowledgments
This paper was supported by the National Natural Science Foundation of China (No. 62306112), Shanghai Sailing Program (No. 23YF1409400), and Shanghai Pilot Program for Basic Research (No. 22TQ1400100-20).
Appendix A Models
Qwen2.5-Instruct series. We choose the {0.5, 1.5, 3, 7, 14, 32, 72} sizes.
Internlm2.5-Chat series. We choose the {1.8, 7, 20} sizes.
Llama3.1-Instruct series. We choose the {8, 70} sizes.
MiniCPM3-4B 888https://huggingface.co/openbmb/MiniCPM3-4B. It’s a lightweight Chinese LLM.
Apollo series We choose the Apollo2 {0.5, 1.5, 7} and Apollo 72. Apollo models trained on Qwen and Qwen2 with a high quality medical dataset.
HuatuoGPT2 series We choose the HuatuoGPT2 {7, 14}. We strictly followed the default load method of HuatuoGPT2-34B, but the inference time was too long, and the final results were not satisfactory. As a result, we did not conduct further experiments on HuatuoGPT2-34B.
Appendix B Training Details
As shown in Figure 5 and Table 6, Qwen2.5-7B-Instruct is an excellent foundation model, considering both performance and time cost. However, for practical purposes, we prefer using a more lightweight model. We trained the model with LLaMA-Factory. 999https://github.com/hiyouga/LLaMA-Factory We use the default ds_z3_config and Lora fine-tuning. Detail parameters: per_device_train_batch_size: 3; gradient_accumulation_steps: 8; learning_rate: 1.0e-5;num_train_epochs:4; lr_scheduler_type: cosine; warmup_ratio: 0.1; fp16: true; ddp_timeout: 180000000.
Appendix C Discussion on Test-Time Scaling
While test-time scaling like (CoT-SC, ToT or O1, R1) has attracted significant research attention, practical deployment in clinical settings requires careful consideration of GPU resource constraints, as hospitals typically have limited computational capacity. To address this, we quantitatively compare the inference efficiency of our method against CoT in Table 6, with measurements reported in GPU-hours.
| zero-shot | one-shot | |||
| CoT | CF-IR | CoT | CF-IR | |
| Qwen2.5-7B | 0.59 | 0.46 | 0.56 | 0.69 |
| Qwen2.5-14B | 1.30 | 1.26 | 1.40 | 1.70 |
| Qwen2.5-32B | 3.88 | 3.60 | 3.32 | 4.16 |
| Qwen2.5-72B | 7.60 | 7.04 | 5.80 | 6.84 |
| Models | CoT | CF-IR |
| DeepSeek-V3 | 3h40min | 3h29min |
| DeepSeek-R1 | 10h10min | 9h30min |
Comparing inference times between chat and reasoning models in Table 7, we observe that while the reasoning model demonstrates better scalability, the performance gains remain marginal in Table 8. In the zero-shot setting, R1 did not demonstrate significant improvement. In contrast, our CF-IR framework outperformed the long cot method.
| Standard | CoT | CF-IR | |
| Qwen2.5-7B | 66.49 | 82.80 | 82.92 |
| Qwen2.5-14B | 78.79 | 86.49 | 87.21 |
| Qwen2.5-32B | 75.54 | 82.03 | 86.11 |
| Qwen2.5-72B | 87.77 | 87.51 | 92.73 |
| llama3.1-70B | 82.54 | 85.85 | 85.47 |
| GPT-4o | 77.45 | 88.91 | 91.84 |
| DeepSeek-V3 | 82.67 | 86.83 | 91.84 |
| DeepSeek-R1 | 81.01 | 82.99 | 92.73 |
| Fileds Name | Explanation |
| definition | The definition of the indicator |
| formula | The calculation formula of the indicator |
| significance | The medical impact of indicator |
| other | The relative knowledge or supplement |
| instruction_standard | The standard prompt for MQCIC with the rule |
| numerator | The numerator of indicator |
| denominator | The denominator of indicator |
| rule | The numerator rule of indicator |
| facts | The templated clinical facts list |
| logical_rules | The logical rules lists. Containing natural and symbolic languages. |
| Fileds Name | Explanation |
| unique_id | The unique id of the indicator |
| patient note | The patient note of the instance |
| explaination | The explanation of the answer |
| label | The label of the answer |
| facts | The list that contains all templated clinical facts with the related original text and answer |
| logic | The list that contains logical rules with the answer |
Appendix D Prompt
Here are the zero-shot prompt templates for the data construction, rule representation enhancement, clinical fact-based inferential rule reasoning method, and other prompt-based methods we used in this paper.
For each indicator rule question, we provide a list of facts that should be extracted. The prompt for clinical fact extraction is shown in Figure 7. The main fields are in the Table 9 and Table 10.
Prompt for Rule Enhancement. As mentioned in Section 4, we leverage LLMs to transform rules through three steps: Knowledge Enhancement, Rule Decomposition, and Clinical Fact Templazation. As shown in Figure 8, 9 and 10, we leverage GPT-4o first to generate the relative knowledge, logical rules, and templated clinical facts separately. Prompt for Different Methods. The detail prompt of different method as shown in Figure 11, Figure 12 and Figure 13.
Appendix E Additional Analysis Details
E.1 Error Analysis Details
As shown in Table 11 and Table 12 we observed: (1) Although the one-shot method reduces errors across both types in both approaches, the CoT method still results in more Type B errors, which may be due to the differing reasoning paths in the examples. (2) The CF-IR method effectively reduces Type B errors, but when it comes to Type A errors, the issue seems to be more related to the model’s intrinsic capabilities, which our method has not been able to enhance or activate effectively.
E.2 Clinical Fact Verification and Inferential Rule
It may be due to the excessively strict scoring criteria that the overall score for clinical fact verification ability is relatively low, but the trend still aligns with expectations. As mentioned in Section 6.3, there is a clear correlation between the model parameters and capabilities in the Apollo and Qwen series. As shown in Table 13, however, models like llama3.1 and HuatuoGPT2, due to differences in the number of parameters, fail to demonstrate this relationship.
| Zero-shot CoT | One-shot CoT | ||||||||
| Type A | Type B | Type C | Total | Type A | Type B | Type C | Total | ||
| General | MiniCPM3-4B | 0.21 | 0.08 | 0.00 | 0.28 | 0.13 | 0.03 | 0.00- | 0.16 |
| Internlm2.5-1.8B | 0.35 | 0.08 | 0.00 | 0.44 | 0.23 | 0.08- | 0.00- | 0.31 | |
| Internlm2.5-7B | 0.15 | 0.12 | 0.00 | 0.27 | 0.13 | 0.03 | 0.00- | 0.16 | |
| Internlm2.5-20B | 0.17 | 0.05 | 0.00 | 0.22 | 0.11 | 0.03 | 0.00- | 0.14 | |
| Qwen2.5-0.5B | 0.34 | 0.10 | 0.00 | 0.44 | 0.29 | 0.10- | 0.00- | 0.39 | |
| Qwen2.5-1.5B | 0.21 | 0.16 | 0.00 | 0.37 | 0.24 | 0.05 | 0.00- | 0.29 | |
| Qwen2.5-3B | 0.16 | 0.11 | 0.00 | 0.27 | 0.15 | 0.07 | 0.00- | 0.22 | |
| Qwen2.5-7B | 0.13 | 0.04 | 0.00 | 0.17 | 0.12 | 0.03 | 0.00- | 0.15 | |
| Qwen2.5-14B | 0.10 | 0.08 | 0.00 | 0.18 | 0.08 | 0.04 | 0.00- | 0.12 | |
| Qwen2.5-32B | 0.09 | 0.05 | 0.00 | 0.14 | 0.06 | 0.04 | 0.00- | 0.10 | |
| Qwen2.5-72B | 0.09 | 0.03 | 0.00 | 0.12 | 0.07 | 0.02 | 0.00- | 0.09 | |
| llama3.1-8B | 0.23 | 0.14 | 0.00 | 0.37 | 0.14 | 0.04 | 0.00- | 0.18 | |
| llama3.1-70B | 0.10 | 0.04 | 0.00 | 0.14 | 0.08 | 0.03 | 0.00- | 0.11 | |
| GPT-4o | 0.08 | 0.03 | 0.00 | 0.11 | 0.06 | 0.02 | 0.00- | 0.08 | |
| Medical | HuatuoGPT2-7B | 0.17 | 0.28 | 0.00 | 0.46 | 0.23 | 0.23 | 0.02 | 0.47 |
| HuatuoGPT2-14B | 0.32 | 0.13 | 0.01 | 0.45 | 0.28 | 0.18 | 0.01- | 0.48 | |
| Apollo2-0.5B | 0.34 | 0.19 | 0.04 | 0.58 | 0.30 | 0.15 | 0.02 | 0.46 | |
| Apollo2-1.5B | 0.23 | 0.25 | 0.00 | 0.48 | 0.27 | 0.07 | 0.00- | 0.34 | |
| Apollo2-7B | 0.17 | 0.23 | 0.00 | 0.40 | 0.24 | 0.04 | 0.00- | 0.28 | |
| Apollo-72B | 0.14 | 0.10 | 0.00 | 0.24 | 0.11 | 0.03 | 0.00- | 0.14 | |
| Zero-shot CF-IR | One-shot CF-IR | ||||||||
| Type A | Type B | Type C | Total | Type A | Type B | Type C | Total | ||
| General | MiniCPM3-4B | 0.27 | 0.04 | 0.00 | 0.31 | 0.11 | 0.06 | 0.00- | 0.17 |
| Internlm2.5-1.8B | 0.39 | 0.07 | 0.00 | 0.46 | 0.27 | 0.10 | 0.00- | 0.36 | |
| Internlm2.5-7B | 0.18 | 0.03 | 0.00 | 0.21 | 0.13 | 0.02 | 0.00- | 0.16 | |
| Internlm2.5-20B | 0.15 | 0.04 | 0.00 | 0.19 | 0.10 | 0.01 | 0.00- | 0.11 | |
| Qwen2.5-0.5B | 0.28 | 0.18 | 0.00 | 0.46 | 0.28- | 0.18- | 0.01 | 0.46 | |
| Qwen2.5-1.5B | 0.36 | 0.02 | 0.00 | 0.38 | 0.25 | 0.02- | 0.00- | 0.27 | |
| Qwen2.5-3B | 0.31 | 0.01 | 0.00 | 0.32 | 0.16 | 0.02 | 0.00- | 0.18 | |
| Qwen2.5-7B | 0.14 | 0.03 | 0.00 | 0.17 | 0.08 | 0.01 | 0.00- | 0.10 | |
| Qwen2.5-14B | 0.12 | 0.01 | 0.00 | 0.13 | 0.07 | 0.01- | 0.00- | 0.08 | |
| Qwen2.5-32B | 0.12 | 0.02 | 0.00 | 0.14 | 0.05 | 0.00 | 0.00- | 0.05 | |
| Qwen2.5-72B | 0.06 | 0.01 | 0.00 | 0.07 | 0.03 | 0.01- | 0.00- | 0.04 | |
| llama3.1-8B | 0.20 | 0.02 | 0.00 | 0.22 | 0.12 | 0.01 | 0.00- | 0.14 | |
| llama3.1-70B | 0.12 | 0.03 | 0.00 | 0.15 | 0.07 | 0.00 | 0.00- | 0.08 | |
| GPT-4o | 0.07 | 0.01 | 0.01 | 0.08 | 0.05 | 0.01- | 0.00 | 0.06 | |
| Medical | HuatuoGPT2-7B | 0.33 | 0.16 | 0.01 | 0.50 | 0.31 | 0.11 | 0.02 | 0.43 |
| HuatuoGPT2-14B | 0.43 | 0.09 | 0.02 | 0.54 | 0.46 | 0.08 | 0.04 | 0.57 | |
| Apollo2-0.5B | 0.32 | 0.15 | 0.12 | 0.59 | 0.27 | 0.07 | 0.01 | 0.35 | |
| Apollo2-1.5B | 0.30 | 0.17 | 0.01 | 0.49 | 0.26 | 0.08 | 0.01- | 0.35 | |
| Apollo2-7B | 0.29 | 0.10 | 0.00 | 0.38 | 0.14 | 0.21 | 0.01- | 0.35 | |
| Apollo-72B | 0.26 | 0.01 | 0.00 | 0.27 | 0.11 | 0.03 | 0.00- | 0.14 | |
| Clinical Fact Verification | Inferential Rule Reasoning | One-Stage | ||||
| Faithfulness | Correctness | NL(ACC) | SY(ACC) | Best(ACC) | ||
| General | MiniCPM3-4B | 45.14 | 36.05 | 67.38 | 69.68 | 72.10* |
| Qwen2.5-0.5B | 45.37 | 40.1 | 74.90 | 75.92 | 56.05* | |
| Qwen2.5-1.5B | 46.45 | 43.67 | 75.66 | 74.64 | 66.11* | |
| Qwen2.5-3B | 52.08 | 50.02 | 83.69 | 81.18 | 76.05* | |
| Qwen2.5-7B | 67.63 | 53.13 | 90.44 | 88.91 | 82.92* | |
| Qwen2.5-14B | 67.94 | 62.35 | 92.35 | 94.52 | 87.21* | |
| Qwen2.5-32B | 75.41 | 69.47 | 89.80 | 94.14 | 86.11* | |
| Qwen2.5-72B | 74.07 | 77.13 | 93.63 | 93.88 | 92.73* | |
| llama3.1-8B | 47.89 | 37.88 | 82.67 | 80.89 | 78.34* | |
| llama3.1-70B | 48.59 | 40.48 | 85.60 | 85.98 | 85.47* | |
| Medical | HuatuoGPT2-7B | 13.27 | 22.93 | 47.89 | 44.20 | 54.26* |
| HuatuoGPT2-14B | 32.55 | 35.47 | 62.42 | 62.03 | 55.28* | |
| Apollo2-0.5B | 5.19 | 10.73 | 9.80 | 18.47 | 49.29* | |
| Apollo2-1.5B | 25.31 | 29.21 | 48.53 | 52.86 | 55.03* | |
| Apollo2-7B | 33.78 | 33.93 | 56.30 | 42.29 | 61.91* | |
| Apollo-72B | 42.31 | 32.07 | 76.81 | 71.59 | 76.24* | |
| Average | 45.18 | 42.16 | 71.12 | 70.69 | 70.94* | |